# TacoGFN: Target-conditioned GFlowNet for Structure-based Drug Design

### v5

@@ 4.3 Docking score predictor with pharmacophore prior
The exploration of chemical space with a GFlowNet requires evaluating binding affinities for millions of molecules sampled during training. However, using molecular docking for evaluation is computationally expensive. Here, we propose a fast ML-based docking score predictor that generalizes well across molecule and protein pocket distributions, as described in Figure 3 and Equation 5. (See details in Appendix C)
We leverage PharmacoNet (Seo & Kim, 2023), a recent deep learning method to obtain the pharmacophore $\mathbb{P}$ from a pocket structure $p$. 111 Pharmacophore is a point set, where each point describes the desirable motif properties (such as being aromatic) a ligand should possess at this geometric position to form energetically or entropically favourable interactions (e.g. $\pi$$\pi$ stacking) with the protein target. (Wermuth et al., 1998; Yang, 2010). To enrich the representation, we obtained the embedding of the pocket $\mathbf{z}^{p}$ and the embeddings of pharmacophore points $\{\hat{\mathbf{h}}^{\mathbb{P}}_{i}\}$ from PharmacoNet. The 1D representation vectors of the pharmacophore $\mathbf{z}^{\mathbb{P}}$ is computed by the global pooling of $\{\hat{\mathbf{h}}^{\mathbb{P}}_{i}\}$
@@ 5 Experiments
5.1 Datasets
We train the docking score predictor on datasets containing protein-ligand pairs and their corresponding Quick Vina 2.1 (QVINA) docking scores. We consider two datasets to train our predictor. The first dataset - CrossDock-100k is commonly used in previous works (Luo et al., 2021; Peng et al., 2022; Guan et al., 2023b). We first apply identical splitting and processing protocol on the CrossDocked dataset (Francoeur et al., 2020) to obtain the same split of 100k protein-ligand pairs (See details in Appendix D.1). We then dock and score using Quick Vina 2.1 (QVINA) (Trott & Olson, 2010; Alhossary et al., 2015) for all protein-ligand pairs from this training split. For our purpose of training a docking score predictor, we do not require high-quality protein-ligand structural data such as the CrossDock-100k set. Thus, we introduce a second, novel dataset for docking score prediction more suited for our method called ZINCDock-15M. It consists of about 15M docking simulation data - from docking 1,000 random ZINC20 (Irwin et al., 2020) molecules into each of the 15,207 unique pockets from CrossDock-100k training split using QVina.
We could directly compare our molecular generation performance with existing methods when trained on CrossDock-100k, as we train on the same set of protein-ligand complexes. However, we could not fairly train existing methods on the new ZINCDock-15M dataset. Instead of containing experimental-like structure data between known binders and their target, ZINCDock-15M contains random molecules docking conformation in protein pockets. Learning a distribution of random molecules with a distribution learning objective is likely unhelpful for generating high-affinity binding molecules. Our method’s advantage here is that we benefit strongly from this large-scale docking data.
5.2 Evaluation settings
@@
For all experiments, each method is tasked to produce 100 unique and valid molecules (ligands) for each of the 100 protein pockets from the CrossDock-100k test set.
Evaluation metrics. The goal of structure-based drug discovery is to propose novel drug candidates. Importantly, a drug candidate must simultaneously have high affinity, desirable drug-like properties, and be easily synthesizable to be considered for further (expensive) experimental validation. To evaluate these molecular properties, the following metrics are commonly used: (1) Docking Score (DS) approximates the binding energy between a generated molecule and a protein pocket, where a lower docking score indicates a higher binding affinity. Molecules are docked using QVina (Trott & Olson, 2010; Alhossary et al., 2015) (See Appendix D.2 for details). (2) QED is a measure of drug-likeness, estimating a molecule’s suitability as an oral drug based on its properties (Bickerton et al., 2012).(3) Synthetic Accessibility (SA) estimates difficulty of synthesizing the molecule (Ertl & Schuffenhauer, 2009). The score is normalized between 0 and 1 using the formula $(10-SA)/9$.
We report the average of these performance metrics in Table 2. However, presenting only the averages of these property values is insufficient, as it ignores whether any individual molecule simultaneously satisfies all of the necessary requirements (Lee et al., 2023). Nor does it evaluate the ability of the model to generate novel structures, which is important for avoiding already patented scaffolds or known binders as noted in (Schreyer & Blundell, 2012).
Therefore, following previous work in de novo drug discovery (Lee et al., 2023), we define a desirable (desr.) molecule candidate for drug as a molecule with a QED > 0.5, and an SA > 0.5. A hit molecule is defined as a molecule with a Docking Score (DS) < -8.0 in addition to satisfying the desirable requirements. Lastly, the novelty (nov.) constraint selects molecules with a maximum Tanimoto similarity < 0.5 with the CrossDocked training molecules, which represent the known binder set. We employ the following additional metrics to better reflect the drug discovery objectives: (4) Hit (%) is the fraction of generated molecules that satisfy the hit constraint. (5) Novel Hit (%), introduced in (Lee et al., 2023), is the fraction of generated molecules that satisfy both the novelty and hit constraint. (6) Top-10 DS is the average of top 10 docking scores of all generated molecules. (7) Top-10 Desr. DS is the average of the top 10 docking scores of generated molecules that satisfy the desirable constraint. (8) Novelty (Lee et al., 2023; Peng et al., 2022) is the average Tanimoto distance of molecules to the nearest ligand from the CrossDocked training set (known binders) (Bajusz et al., 2015). (9) Diversity is calculated as the average pairwise fingerprint Tanimoto distance between molecules generated for a pocket. (9) Time is the average runtime (in seconds) for generating 100 unique and valid molecules for a pocket.
@@
Table 1: Evaluation of 100 generated molecules for each of 100 protein pockets from the CrossDocked test set. The best results are in bold and the second best results are underlined. The average and median values are calculated over the averages for each pocket.
Novel Top-10 DS ($\downarrow$) Top-10 desr. DS ($\downarrow$) Nov. ($\uparrow$) Div. ($\uparrow$) Time ($\downarrow$)
Model Hit % ($\uparrow$) Hit % ($\uparrow$) Avg. Med. Avg. Med. Avg. Med. Avg. Med.
Random ZINC 28.88 17.86 -9.12 -9.23 -9.05 -9.16 0.55 0.55 0.76 0.76 -
Pocket2Mol (Peng et al., 2022) 20.85 6.79 -9.10 -8.4 -8.47 -8.23 0.45 0.46 0.74 0.77 2503
TargetDiff (Guan et al., 2023a) 15.65 7.94 -9.40 -9.54 -8.52 -8.70 0.49 0.49 0.74 0.73 3428
DecompDiff (Guan et al., 2023b) 8.95 3.01 -9.79 -9.69 -4.92 -5.28 0.45 0.45 0.60 0.60 6189
TacoGFN (ours) 58.25 43.49 -9.97 -10.26 -9.96 -10.26 0.52 0.52 0.53 0.53 4.29
[…]
Table 2: Comparison of the properties of the reference molecules and the generated molecules for the CrossDocked test set pockets. The reference molecules are from the CrossDocked test set.
[…]
Baselines. We compare TacoGFN against several state-of-the-art SBDD methods. In particular, we select three recent methods - Pocket2Mol (Peng et al., 2022), TargetDiff (Guan et al., 2023a), DecompDiff (Guan et al., 2023b) and reproduce their experimental results (See Appendix D.4 for details). Except for the results of these three methods, all other evaluation results are taken from the recent SBDD survey paper (Zhang et al., 2023c). Furthermore, we randomly select 100 molecules from ZINC20 to form our baseline Random ZINC.
[…]
Figure 4: The average of the top-10 docking score of molecules generated for individual CrossDocked test pockets (target) by DecompDiff and TacoGFN. Targets are sorted by the average of the top-10 docking score of TacoGFN generated molecules. A lower docking score means a higher estimated binding affinity. Color is used to denote the average QED value of molecules in the Top-10 set. A higher QED indicates the molecule is more drug-like.
@@ 5.3 Experimental results
Table 1 and Table 2 present the aggregate of metrics highly important for the objective of de novo structure-based drug design. We examine the docking score performances for individual pockets in Figure 4 and show examples of generated molecules in appendix Figure 5. Lastly, we conduct ablation studies to show the benefits of using a larger docking score dataset in Table 3 and validate the utility of the pocket conditioning in Table 4.
De novo drug design. Table 1 shows that TacoGFN achieves an impressive hit rate of $58.25\%$, a large improvement compared to Pocket2Mol $(20.85\%)$. Notably, TacoGFN achieves the best average for Top-10 DS $(-9.97)$. It also has a significantly improved Top-10 desirable DS $(-9.96)$ compared to the second best method TargetDiff $(-8.52)$. Since the vast majority of molecules generated by TacoGFN have good QED and SA values, we observe only a tiny decrease in Top-10 DS average when we filter for desirable molecules only. In contrast, the Top-10 DS average noticeably drops for all baseline methods if we consider only desirable molecules. Finally, we note that TacoGFN is the only method that consistently outperforms the Random ZINC baseline in terms of all affinity-focused metrics, i.e. except novelty and diversity.
Furthermore, TacoGFN achieves an impressive $5$-fold improvement in the novel hit rate compared to the state-of-the-art methods. It also achieves the best novelty amongst the three selected representative SBDD methods. These results demonstrate our approach is able to search the greater chemical space and discover highly novel hit molecules, instead of memorizing the training set.
In Table 2, we show that TacoGFN and its variant outperform all baselines in terms of average and median of docking score, QED and SA. This confirms that TacoGFN can generate high-affinity molecules that simultaneously satisfy the drug-likeness and ease-of-synthesis requirements.
As reported in Table 1, the molecules generated by TacoGFN for one particular pocket have lower diversity than those generated by the distribution learning methods that do not attempt to satisfy multiple objectives. As TacoGFN seeks to generate molecules that satisfy multiple property requirements and bind strongly to a specific pocket structure, the resulting solution set is inevitably smaller, resulting in reduced diversity.
Figure 4 shows that TacoGFN achieves better average Top-10 docking scores than DecompDiff in $58\%$ of the test pockets. Notably, the top molecules generated by TacoGFN consistently demonstrate high QED values, whereas the top molecules generated by DecompDiff are characterized by a low QED value. This figure shows that TacoGFN achieves superior docking scores without compromising the drug-like properties across all pocket structures. In appendix Figure 5, we show TacoGFN is a able to generate molecules with significantly improved docking scores compared to native ligands.
@@
Generation runtime and quality. The generation process of TacoGFN is a few orders of magnitude faster than existing autoregressive or diffusion-based generative methods. Additionally, TacoGFN achieves 100% in validity and uniqueness, demonstrating the efficiency of our generation framework. As shown in Table 8 of Appendix D.5.2, this is a significant improvement over previous methods which often generate invalid or duplicate molecules.
[…]

### v6

@@ 4.3 Docking score predictor with pharmacophore prior
The exploration of chemical space with a GFlowNet requires evaluating binding affinities for millions of molecules sampled during training. However, using molecular docking for evaluation is computationally expensive. Here, we propose a fast ML-based docking score predictor that generalizes well across molecule and protein pocket distributions, as described in Figure 2 and Equation 5. (See details in Appendix C.2)
We leverage PharmacoNet (Seo & Kim, 2023), a recent deep learning method to obtain the pharmacophore $\mathbb{P}$333 Pharmacophore is a point set, where each point describes the desirable motif properties (such as being aromatic) a ligand should possess at this geometric position to form energetically or entropically favourable interactions (e.g. $\pi$$\pi$ stacking) with the protein target. (Wermuth et al., 1998; Yang, 2010). from a pocket structure $P$.To enrich the representation, we obtained the embedding of the pocket $\mathbf{z}^{P}$ and the embeddings of pharmacophore points $\{\hat{\mathbf{h}}^{\mathbb{P}}_{i}\}$ from PharmacoNet. The 1D representation vectors of the pharmacophore $\mathbf{z}^{\mathbb{P}}$ is computed by the global pooling of $\{\hat{\mathbf{h}}^{\mathbb{P}}_{i}\}$
@@ 5 Experiments
5.1 Dataset
We train and evaluate TacoGFN on the commonly used CrossDocked benchmark (Francoeur et al., 2020). We first apply the splitting and processing protocol on the CrossDocked dataset to obtain the same train and test split of the 100k protein-ligand pairs as previous methods (Luo et al., 2021; Peng et al., 2022; Guan et al., 2023b). (See details in Appendix D.1). We then train our docking score predictor on the training split of the protein-ligand pairs to predict their corresponding Vina Dock scores. Since TacoGFN is trained on the same set of protein-ligand pairs and evaluated on the same unseen pockets, our experimental results can be directly compared against the published results of existing structure-based generative models.
5.2 Training
We train one TacoGFN model to generate molecules conditional to any protein target structures using predicted affinity, Synthetic Accessibility (SA), and drug-likeness (QED) as the reward. We also adopt Double GFN (Lau et al., 2023) to improve exploration in sparse reward domains and high-dimensional states, by initializing two networks to model the action policy: an online network and a target network.
For each training trajectory, one protein pocket is randomly drawn from the CrossDocked training set first. Then, a molecule is generated through sampling a sequence of actions using our target network which models the forward action policy $P_{target}$. The reward is then computed for that molecule with respect to the protein target. The Trajectory Balance loss (Equation 1) is used to train the online network modelling the action policy $P_{online}$. This loss has the objective for our model to generate objects with probabilities proportional to the rewards (consisting of predicted docking score, QED and SA). Periodically, the weights of the target network are updated by the weights of the online network using a delayed strategy. This strategy reduces training instability and promotes explorations of the larger chemical space.
5.3 Evaluation
@@
In all evaluations, each structure-based generative model is tasked to produce 100 molecules (ligands) for each of the 100 unseen protein pockets from the CrossDock-100k test set.
Evaluation metrics. We adopt the following commonly used metrics from Guan et al. (2023a) and Reidenbach (2024): (1) Validity is the percentage of unique generated molecules free of reconstruction errors and disconnections as determined by RDKit. (2) Vina Dock approximates the binding energy between a generated molecule and a protein pocket, where a lower docking score indicates a higher binding affinity. (3) High Affinity measures the percentage of generated molecules with higher affinity than the reference molecule. (4) QED is a measure of drug-likeness, estimating a molecule’s suitability as an oral drug based on its properties (Bickerton et al., 2012). (5) Synthetic Accessibility (SA) estimates difficulty of synthesizing the molecule (Ertl & Schuffenhauer, 2009). The score is normalized between 0 and 1 using the formula $(10-SA)/9$. (6) Diversity is calculated as the average pairwise fingerprint Tanimoto distance between molecules generated for a pocket. (7) Success Rate is the percentage of molecules which pass the same criteria (QED > 0.25, SA > 0.59, Vina Dock < -8.18) as in Long et al. (2022); Guan et al. (2023b); Zhou et al. (2024); Reidenbach (2024). (8) Time is the average runtime (in seconds) for generating 100 unique and valid molecules for a pocket.
Baselines and problem settings. Similar to Zhou et al. (2024) and (Reidenbach, 2024), we separate existing methods by their problem definitions:
1) Generative methods are expected to generalize for pocket structures unseen during training. They generate molecules for test pockets at inference time without optimization loops or access to docking programs. Therefore, methods under this setting are only allowed to generate 100 molecules for each pocket in one-shot. We compare TacoGFN trained on CrossDocked-100k against the following generative models: liGAN (Ragoza et al., 2022), GraphBP (Liu et al., 2022), AR (Luo et al., 2021), Pocket2Mol (Peng et al., 2022), TargetDiff (Guan et al., 2023a), DiffSBDD (Schneuing et al., 2023), DecompDiff (Guan et al., 2023b). For consistency with existing baselines, molecules are docked using QVina in this problem setting (Trott & Olson, 2010; Alhossary et al., 2015). For details on the docking protocol for the generative setting, please see Appendix D.2.
2) Optimization methods are able to leverage docking on the target pocket. Unlike the generative setting, these methods typically iteratively optimize the candidate pool and select the top 100 molecules. We compare against the following methods: RGA (Fu et al., 2022) and EvoSBDD (Reidenbach, 2024) use a black-box algorithms to conduct rounds of the optimization process, with molecule fitness based on docking score to the target pocket. DecompOpt and TargetDiff+Opt (Zhou et al., 2024) optimizes molecules generated by pre-trained SBDD models via rounds of optimization process involving re-docking to the target pocket. TacoGFN+FT fine-tunes a pre-trained TacoGFN to tailor the model to the target pocket using docking as the reward.
@@
Table 1: Comparison of the properties of molecules generated in the generative problem setting for the CrossDocked test set pockets. The reference molecules are from the CrossDocked test set. The best results are in bold. The average and median values are calculated over the averages for each pocket. The prior method results are taken from their publication.
Validity ($\uparrow$) Vina Dock ($\downarrow$) High Affinity ($\uparrow$) QED ($\uparrow$) SA ($\uparrow$) Diversity ($\uparrow$) Success Rate ($\uparrow$) Time ($\downarrow$) Model Avg. Avg. Med. Avg. Med. Avg. Med. Avg. Med. Avg. Med. Avg. Gen Reference - -7.45 -7.26 - - 0.48 0.47 0.73 0.74 - - 25.0% - liGAN - -6.33 -6.20 21.1% 11.1% 0.39 0.39 0.59 0.57 0.66 0.67 3.9% - GraphBP - -4.80 -4.70 14.2% 6.7% 0.43 0.45 0.49 0.48 0.79 0.78 0.1% 10 AR 92.95% -6.75 -6.62 37.9% 31.0% 0.51 0.50 0.63 0.63 0.70 0.70 7.1% 19659 Pocket2Mol 98.31% -7.15 -6.79 48.4% 51.0% 0.56 0.57 0.74 0.75 0.69 0.71 24.4% 2504 TargetDiff 90.35% -7.80 -7.91 58.1% 59.1% 0.48 0.48 0.58 0.58 0.72 0.71 10.5% 3428 DiffSBDD 85.01% -8.03 -7.75 55.3% 56.6% 0.47 0.47 0.55 0.56 0.76 0.76 6.0% 160 DecompDiff 71.96% -8.39 -8.43 64.4% 71.0% 0.45 0.43 0.61 0.60 0.68 0.68 24.5% 6189 TacoGFN (Ours) 100% -8.24 -8.44 67.5% 92.0% 0.67 0.67 0.79 0.79 0.53 0.53 56.0% 4
@@ 5.4 Experimental results
Table 1 compares TacoGFN against other baselines in the generative problem setting for SBDD. We examine the docking score performances for individual pockets in Figure 4 and 5. We also show examples of generated molecules in Figure 3 and analyze their average physical properties in Table 2. We then compare TacoGFN+FT against existing methods for the optimization problem setting in Table 3 and 4, where docking is used in optimization rounds. Lastly, we conduct ablation studies to show the benefits of using a larger docking score dataset in Table 5 and validate the utility of the pocket conditioning in Table 6.
Generative setting. Table 1 highlights the strong performance of TacoGFN in the generative problem setting for the CrossDocked test set pockets. Notably, TacoGFN boasts a significant improvement in success rate at 56.0%, more than doubling the previous best of 24.5% achieved by DecompDiff. This improvement is due to TacoGFN’s ability to generate high-affinity molecules that simultaneously satisfy the drug-likeness and ease-of-synthesis requirements. In fact, TacoGFN records the best QED, SA, and High Affinity simultaneously among all generative methods. In Figure 3, we show examples of molecules generated by TacoGFN and show its ability to generate molecules with significantly improved docking scores compared to native ligands.
It is difficult to discover molecules that simultaneously exhibit better QED and Vina Dock compared to known binders and molecules from existing baselines due to the trade-off between Vina Dock and QED. Molecules with higher molecular weight are more likely to have strong Vina Dock because of the presence of more interacting atoms, but they tend to have worse drug-like properties (QED).  [^4] This trade-off is clearly observed with Pocket2Mol and DecompDiff (see Table 1 and 2), where DecompDiff has strong Vina Dock but lower QED, while Pocket2Mol exhibits the opposite trend. In Figure 4, we show the performance breakdown for the 100 test protein pockets. TacoGFN achieves better average Top-10 Vina Dock than DecompDiff in $57\%$ of the test pockets. Notably, the top molecules generated by TacoGFN also consistently demonstrate higher QED values. As shown in Appendix Figure 5, in the pockets where DecompDiff achieves a lower Top-10 Vina Dock, the molecules often exhibit a molecular mass larger than 500 daltons. These heavier molecules violate the ideal properties of orally active drugs according to the Rule of 5 (Lipinski et al., 1997). In addition, heavy molecules with high docking scores are more likely to be false positives (Pan et al., 2002). TacoGFN is able to find molecular spaces that improve both QED and Vina Dock requirements not only by modelling the reward distribution but also by exploring a broader chemical space. This is achieved through learning from generated examples using our online policy, rather than being limited to the training data.
@@
We note exploring millions of molecules online could not be easily achieved with existing SBDD baseline. This is because the generation process of TacoGFN is a few orders of magnitude faster than existing autoregressive or diffusion-based generative methods. Additionally, TacoGFN achieves 100% in validity and uniqueness, demonstrating the efficiency of our fragment-based 2D generation framework. As shown in Table 1, TacoGFN achieves a significant improvement in both time and validity over previous methods.
Molecules generated from TacoGFN have more ideal molecular weight and drug-likeness properties, in addition to achieving better Vina Dock; Therefore, they are more suitable as drug candidates. TacoGFN samples at reward temperature $\beta$ of 64 at inference time, resulting in the modelling of the probability $p(x|c)\propto R(x,c)^{64}$. This policy focuses more on the modes of reward function. Therefore the molecules sampled have higher quality but lower diversity than those generated by the methods that do not attempt to satisfy multiple objectives. By changing $\beta$ or reward function we can trade off rewards with diversity as shown in Appendix D.4.2.
Figure 3: Our method is focused on de novo hit discovery - finding novel and diverse high-scoring hits for a protein target. Our method does not optimize based on a given reference seed compound. The goal of reward-based sampling is to sample diverse high-scoring molecules. To provide a fair overview of the model’s performance, we selected protein pockets 4iwq and 4q8b, which are at the 25th and 75th percentiles, respectively, based on their docking scores with their native ligands. We show compare the molecules generated by TacoGFN against the native ligand with their QED, SA, Novelty, and Docking score (Vina).
Figure 4: The average of the top-10 Vina Dock of molecules generated for individual CrossDocked test pockets (target) by DecompDiff and TacoGFN. Targets are sorted by the average of the top-10 docking score of TacoGFN generated molecules. A lower docking score means a higher estimated binding affinity. Color is used to denote the average QED value of molecules in the Top-10 set. A higher QED indicates the molecule is more drug-like.
Evaluation of physical properties.
We evaluate the following physical properties important for small molecule drugs in Table 2: (1) Ideal Mol. Wt. is the percentage of molecules satisfying molecular weight within 160 - 480 daltons - the typical acceptable molecular weight range for small molecule drugs (Ghose et al., 1999). (2) Mol. Wt. is the molecular weight of molecules in Daltons. (3) Num Heavy Atom is the heavy atom (any atom that is not hydrogen) count of the molecules generated. (4) Strain Energy is the difference between the re-docked poses (meaning generated molecules are docked with RDKit’s ETKDG conformation initialization), versus the strain energy of their unstrained ETKDG poses. We consider the sum of torsional, steric and electrostatic strains. If the strain energy of the ligand’s docked pose is very high, the physical compatibility for such binding becomes poor (Perola & Charifson, 2004; Gu et al., 2021).
For baseline, we compare against the top three generative methods ranked by success rate: DecompDiff Guan et al. (2023b), Pocket2Mol (Peng et al., 2022) and TargetDiff Guan et al. (2023a). Additionally, we sample a random set of molecules from the ZINC virtual library for comparison.
Table 2: Comparison on the number of heavy atoms, strain energies
Ideal Mol. Wt. ($\uparrow$) Mol. Wt. Num Heavy Atom Strain Energy ($\downarrow$)
Model Avg. Avg. Med. Avg. Med. Avg. Med.
ZINC molecules 98.73% 337.63 337.34 23.53 23.50 377.77 375.17
Pocket2Mol 64.55% 248.65 230.99 18.29 16.86 254.21 218.52
TargetDiff 78.37% 323.64 327.31 22.78 23.15 600.31 547.86
DecompDiff 51.18% 494.05 487.51 34.93 34.00 834.72 781.34
TacoGFN (ours) 99.76% 402.56 402.65 30.47 30.52 360.51 357.33
As shown in Table 2, 99.76% of molecules generated by TacoGFN have molecular weights within the 160 to 480 dalton range - the ideal range for small molecule drugs. This ratio is much higher than those of existing methods such as DecompDiff (51.18%), Pocket2Mol (64.55%) and TargetDiff (78.37%). For Pocket2Mol, molecules fall outside of the ideal molecular weight range because they are too small (below 160 Daltons); While for DecompDiff, most molecules are outside of range because they are too heavy (greater than 480 Daltons). We note there are indeed drug candidates with molecular weights or drug-likeness outside of the typical acceptable range. However, they may be regarded as an exception rather than the norm (from a specialized drug class) or have elevated risks for poor absorption and bioavailability (Ritchie & Macdonald, 2014).
PDBbind (Liu et al., 2014), the parent database of the CrossDocked set (Francoeur et al., 2020), contains a large number of common biomolecules such as ATP (molecular weight: 508 daltons) or sulphate ion (molecular weight: 96 daltons). This inherent bias in the training dataset hinders data distribution learning-based generative models from generating molecules that would be acceptable as real drugs. In contrast, TacoGFN learns the actual physical property distribution of drugs, resulting in an average molecular weight (402.56 daltons) of the generated molecules that closely matches that of FDA-approved drugs ($\sim$ 430 daltons).
With regards to strain energy, TacoGFN attains far lower strain energy than DecompDiff and TargetDiff. The strain energy of TacoGFN is comparable with that of ZINC molecules, as shown in Table 2. Pocket2Mol achieves lower strain than ZINC and all other methods, due to the small molecular weight they generate. In summary, this analysis supports that molecules generated by TacoGFN show appropriate physical properties and compatibility with its binding pocket.
Table 3: Comparison of the properties of molecules generated in the optimization problem setting for the CrossDocked test set pockets. The prior method results are taken from their publication. The number of fine-tuning steps is 300 unless otherwise specified.
Validity ($\uparrow$) Vina Dock ($\downarrow$) High Affinity ($\uparrow$) QED ($\uparrow$) SA ($\uparrow$) Diversity ($\uparrow$) Success Rate ($\uparrow$) Time ($\downarrow$)
Model Avg. Avg. Med. Avg. Med. Avg. Med. Avg. Med. Avg. Med. Avg. Gen + Score.
Reference 100% -7.45 -7.26 - - 0.48 0.47 0.73 0.74 - - 25.0% -
RGA - -8.01 -8.17 64.4% 89.3% 0.57 0.57 0.71 0.73 0.41 0.41 46.2% -
TargetDiff+Opt - -8.30 -8.15 71.5% 95.9% 0.66 0.68 0.68 0.67 0.31 0.30 25.8% >3728
DecompOpt - -8.98 -9.01 73.5% 93.3% 0.48 0.45 0.65 0.65 0.60 0.61 52.5% 9241
EvoSBDD ($\alpha$ = 1.3, 140R) 100% -10.27 -10.36 96.5% 100% 0.53 0.52 0.75 0.77 0.63 0.63 78.8% 6300
EvoSBDD ($\alpha$ = 0, $\sigma$ = 1, 140R) 100% -10.14 -10.27 94.4% 100% 0.59 0.59 0.77 0.77 0.62 0.62 86.4% 6300
TacoGFN+FT ($t_{QED}=0.4,t_{SA}=0.6$) 100% -10.78 -10.93 97.1% 100% 0.47 0.47 0.70 0.69 0.62 0.62 87.8% 7750
TacoGFN+FT ($t_{QED}=0.5,t_{SA}=0.8$) 100% -10.33 -10.39 97.0% 100% 0.58 0.58 0.82 0.82 0.61 0.61 88.8% 6230
TacoGFN+FT ($t_{QED}=0.6,t_{SA}=0.8$) 100% -10.14 -10.21 96.2% 100% 0.65 0.65 0.82 0.82 0.59 0.59 86.3% 6140
Optimization setting.
Here we study the optimization problem setting, where methods can conduct optimization rounds leveraging docking program on the target pocket. We evaluate TacoGFN+FT, where we fine-tune a pre-trained TacoGFN, using UniDock (Yu et al., 2023)  [^5] Uni-Dock is a recently proposed GPU-accelerated molecular docking program that achieves more than 2000-fold speed-up compared with the AutoDock Vina running in single CPU core. on the target pocket as the affinity reward for generated molecules. For each protein pocket, we chose to finetune TacoGFN for 300 steps (unless otherwise specified) to match the time for TacoGFN+FT ($t_{QED}=0.5,t_{SA}=0.8$) with EvoSBDD’s (Reidenbach, 2024). We report the time taken for fine-tuning and docking molecules for our method, measured on a single A4000 GPU - a GPU with less performance than the one used by EvoSBDD (A6000 GPU), or TargetDiff+Opt/DecompOpt (A100 GPU). The details of the fine-tuning process and UniDock can be found in Appendix E.
TacoGFN+FT is able to simultaneously optimize a set of potentially conflicting objectives - consisting of Vina Dock, QED and SA. The TacoGFN+FT with balanced objective ($t_{QED}=0.5,t_{SA}=0.8$) simultaneously achieves state-of-art results in Vina Dock, high affinity, synthetic accessibility and success Rate. Since TacoGFN samples molecules during fine-tuning with a lower reward temperature $\beta$ (sampled uniformly between 0 and 64), the diversity of TacoGFN+FT is higher than TacoGFN and comparable with EvoSBDD. Our method variant ($t_{QED}=0.4,t_{SA}=0.6$) focusing on affinity attains the best average Vina Dock of -10.78. Our method variant focusing on QED achieves a comparable QED to the best baseline TargetDiff+Opt (0.65 vs 0.66), but a much better Vina Dock (-10.14 vs -8.30) and SA (0.82 vs 0.68). Please see Appendix Table 12 for full results.
Overall, our proposed fine-tuning pipeline is highly efficient, as TacoGFN+FT surpasses state-of-art performance while taking less time. In comparison to EvoSBDD (Reidenbach, 2024), which does not take pocket structure as input, TacoGFN is already trained to generate drug-like and synthesizable molecules with binding conditions to protein pocket structure. Therefore, finetuning TacoGFN generates a molecular set with a better success rate and high-affinity rate for a new protein pocket. Compared to DecompOpt (Zhou et al., 2024), our method does not need to traverse the 500-1000 denoising steps to generate a molecule. This means we could search for more molecules and discover higher reward candidates in a shorter amount of time.
Table 4 compares the performance of TacoGFN+FT by simply generating an equal number of molecules with base TacoGFN and ranking the top 100 molecules without finetuning (TacoGFN+Rank); Here we show, although TacoGFN+Rank already attains competitive results, our proposed fine-tuning meaningfully further improves Vina Dock and Success Rate.
Table 4: Comparison of performance between TacoGFN with Finetuning (TacoGFN+FT) with simply ranking the same number of generated molecules from TacoGFN by their reward (TacoGFN+Rank).
Validity ($\uparrow$) Vina Dock ($\downarrow$) High Affinity ($\uparrow$) QED ($\uparrow$) SA ($\uparrow$) Diversity ($\uparrow$) Success Rate ($\uparrow$) Time ($\downarrow$)
Model Avg. Avg. Med. Avg. Med. Avg. Med. Avg. Med. Avg. Med. Avg. Gen + Score.
TacoGFN+FT ($t_{QED}=0.5,t_{SA}=0.8$) 100% -10.33 -10.39 97.0% 100% 0.58 0.58 0.82 0.82 0.61 0.61 88.8% 6160
TacoGFN+Rank 100% -10.09 -10.10 97.0% 100% 0.59 0.59 0.82 0.82 0.60 0.61 86.4% 5133